# Hallucination is a Consequence of Space-Optimality: A Rate-Distortion Theorem for Membership Testing

### v1

@@ 1 Introduction
Despite their transformative impact, Large Language Models (LLMs) are hindered by hallucinations – the generation of confident and plausible yet factually incorrect statements. Recently, an influential line of work Kalai et al. (2025); Mohsin et al. (2025); Xu (2025); Wu et al. (2025); Gumaan (2025) theoretically explained the hallucination phenomenon from a statistical learning perspective, considering an LLM as an implementation of a binary classifier on plausible-looking random facts – facts that are unstructured and cannot be logically inferred unless encountered during training, such as phone numbers or biographical details. By a no-free-lunch-style logic, since generalization is impossible on these facts, the model must make uninformed guesses on unseen random facts, leading to systematic inference errors. In the classification setting, hallucination corresponds to false positives resulting from such guesses.
As a natural workaround, researchers have promoted abstention Kalai et al. (2025); Wen et al. (2025) – an LLM is expected to provide an “I don’t know” response on unseen random factual queries. Theoretically, we can further simplify the task by adopting a “closed-world assumption,” where all unseen potential facts are categorized as non-facts for evaluation  [^1] Alternatively, one can assume all true facts are seen during training.. Under this setting, an ideal LLM can be both informative and non-hallucinating on factual queries, as long as it accurately distinguishes the finitely many known facts from all other (potentially true but unknown) non-facts. This is a sample-efficient memorization task.
However, empirical evidence suggests that LLMs struggle to execute this memorization task robustly. Even when permitted to abstain, models continue to generate high-confidence hallucinations, while simultaneously exhibiting “over-refusal” of legitimate queries Cheng et al. (2024); Brahman et al. (2024); Zhu et al. (2025), exhibiting a precision–recall tradeoff. This points to a different bottleneck beyond no-free-lunch: the finite capacity of the model weights forces lossy compression of the training data, making memorization of specific, unstructured facts a non-trivial task.
This information-theoretic perspective connects to several recent works Pan et al. (2025); Shi et al. (2025); Mohsin et al. (2025); Kim (2025b, a) which attribute hallucination to the distortion occurring when compressing an infinitely complex world of knowledge to a finite model capacity. Despite these insights, two gaps remain. First, “compression causes errors” does not explain the shape of the errors: while forgetting is natural under limited capacity, it is unclear why hallucination is particularly prevalent. Second, existing compression-based explanations are either high-level informal arguments or assume an infinite number of facts, while our closed-world assumption restricts facts to a finite set. This motivates a more rigorous study in a simplified setting:
@@
In this work, we argue that these hallucinations are driven by the asymmetry between facts and non-facts. Specifically, if facts are sparse and randomly distributed in a vast universe of plausible statements, then an LLM with limited memory capacity would tend to accept false statements as facts.
To formalize this intuition, we model factuality judgment as a membership testing problem. We view each plausible potential fact as an element $i$ in a universe $\mathcal{U}$, and the set of known facts as a key set $\mathcal{K}\subseteq\mathcal{U}$. Given a query $i\in\mathcal{U}$, the model outputs a confidence score $\hat{x}_{i}\in[0,1]$ indicating its belief that $i\in\mathcal{K}$; under this lens, hallucinations correspond to non-keys receiving high confidence. Query output $\hat{x}_{i}$ can be evaluated under generic error metrics measuring the discrepancy between $\hat{x}_{i}$ and the true label $\mathbbm{1}\{i\in\mathcal{K}\}$. The goal of a membership tester is to use a small memory budget for storing the keys, and achieve small expected error on queries. This abstraction subsumes approximate membership data structures [^2] We will call these data structures “filters” for convenience. such as Bloom filters Bloom (1970), for which the query output is limited to $\{0,1\}$.
@@ 1.1 Our contributions
A rate-distortion theorem for membership testing. We consider the sparse limit where $\frac{|\mathcal{K}|}{|\mathcal{U}|}\to 0$. In this regime, we identify the exact memory–error tradeoff in the form of a rate-distortion theorem. Specifically, the minimum memory budget per key is determined by the minimum KL-divergence between key and non-key output distributions.
@@
Hallucination as the optimal mode of error. Our theory provides two perspectives on why it is memory-efficient for LLMs to hallucinate on random facts. First, consider a model tasked to estimate the probability $\mathbb{P}[i\in\mathcal{K}]$ of a potential fact $i\in\mathcal{U}$ being true, where the output $\hat{x}_{i}\in[0,1]$ is evaluated with logarithmic or cross-entropy loss, consistent with the maximum likelihood objective and previously proposed training/evaluation methods for factual knowledge Kadavath et al. (2022); Cheng et al. (2024).
The unique loss-minimizing query output under this metric, given a fixed memory capacity, is to assign high confidence to all facts, while assigning either zero or fact-level confidence to non-facts. We also empirically verify this tendency on synthetic data. For a probability estimation task, hallucination, instead of systematic forgetting or uniform uncertainty, is the natural mode of error.
Classifier via thresholding and two-sided filters. Our second perspective establishes that any LLM decision mechanism based on score thresholding - whether the scores are derived from generative probabilities Kalai et al. (2025) or the probability estimation output above - is subject to the fundamental memory-error trade-off of two-sided filters, a generalization of Bloom-type filters Bloom (1970) allowing both false positives and false negatives. Unlike the first perspective, this result applies to all LLM-based classification mechanisms and does not assume optimality.
@@
Further discussions on the memory capacity. Although modern models have billions of parameters, the effective memory budget for storing a particular family of random facts can be far smaller. Appendix A provides two explanations: (i) modern networks are encouraged to minimize pure memorization (through various regularization and MDL/PAC-Bayes viewpoints), and (ii) structured knowledge (e.g., language and reasoning) and random facts compete for a finite memory budget, where the former takes precedence during learning. We argue that this implicit bias results in a small effective memory budget for random facts, leading to the observed high-confidence hallucinations.
Conversely, this analysis supports the effectiveness of additional fine-tuning on unstructured random facts, which encourages the model to allocate more memory budget. Our framework also justifies why incorporating external information, such as using RAG Lewis et al. (2020), is effective in mitigating hallucinations: memory budget is no longer a limiting factor when non-parametric memory is present.
@@ 3 A Rate-Distortion Theorem for Membership Testers
Minimizing $F_{p}$ over the feasible sets (distributions $\mu_{K},\mu_{N}$ satisfying error constratints) defines the “rate–distortion” function $R_{p}(\varepsilon_{K},\varepsilon_{N})$. We show that, for each $p$, the “rate” $R_{p}$ is both necessary (Corollary˜3.6) and sufficient (Lemma˜3.7) for achieving the given error level.
@@ 4 Hallucination on Random Facts
Fix a universe $\mathcal{U}$ of $u$ unstructured potential facts; each $i\in\mathcal{U}$ is a plausible natural-language claim (e.g., detailed biographies as in Allen-Zhu and Li (2024)). The set of known facts is the key set $\mathcal{K}\subseteq\mathcal{U}$. In the random-facts regime that motivates our study, $\mathcal{K}$ is a uniformly random subset of size $n$ of $\mathcal{U}$. This regime intentionally singles out non-generalizable factual knowledge, forcing the LLM to behave like a membership tester for $\mathcal{K}$.
Consider an LLM trained to (among other tasks) recognize facts $\mathcal{K}$ from all other non-facts. Viewing the LLM as a membership tester for $\mathcal{K}$, Init corresponds to training on labeled data, and Query corresponds to the inference-time response (or internal evaluation) on a plausible claim $i\in\mathcal{U}$. We study two natural regimes of factuality judgment:
@@
    Probability estimation. The LLM generates a confidence score $\hat{x}_{i}\in[0,1]$ as an estimation of $\mathbb{P}[i\in\mathcal{K}]$.
@@ 4.1 Probability estimation variant
Figure˜1 illustrates the shape of optimal non-fact distribution in blue. The proof appears in Appendix˜C. To solve this optimization problem over probability distributions, we apply variational calculus and check for KKT conditions.
Hallucination channel and its cost. The optimal strategy for a given memory budget is to output a single high confidence $x^{*}$ on all facts, while assigning the same high confidence $x^{*}$ to a $q^{*}$ fraction of non-facts, forming a “hallucination channel.” These non-facts are indistinguishable from facts by any downstream procedure that observes only $\hat{x}$. Moreover, the per-key memory lower bound simplifies to
@@ 4.2 Binary decision variant
We now consider any binary decision on the potential fact universe induced by the LLM. This covers the thresholding of probability estimation in Section˜4.1, or any classifier (e.g., in Kalai et al. (2025)) induced by generative hallucination probabilities. Formally, any such pipeline induces a (randomized) decision $\hat{x}_{i}\in\{0,1\}$ for each $i\in\mathcal{U},$ indicating either acceptance or rejection. We place error constraints on false negative rates (FNR) and false positive rates (FPR):
@@
This memory-error frontier formalizes the tradeoff between hallucination and over-refusal observed in practice Cheng et al. (2024); Brahman et al. (2024). Theorem 4.3 is agnostic to how $\hat{X}$ is produced. Whether the decision is derived from a calibrated probability, a raw logit, a log-likelihood ratio, or a multi-stage verification chain that outputs a scalar score, the final thresholded mapping is exactly a two-sided filter with some $(\varepsilon_{K},\varepsilon_{N})$. Consequently, no choice of threshold can beat the memory-error frontier: a conservative approach to factual answers may reduce $\varepsilon_{N}$, but then either the memory budget or $\varepsilon_{K}$ must increase (more forgetting / abstention).
@@
To complete the argument for filters, in Appendix D.3 we give an explicit construction of a hash-based two-sided filter that achieves the space lower bound up to $o(1)$ bits per key.
@@ Appendix A Detailed Discussion of the Information Budget for Random Facts
We provide a self-contained argument supporting the modeling choice that the relevant memory budget allocated to random facts is orders of magnitude below the raw parameter count.
@@ A.1 Setup: random facts, training data, and learned weights
We consider the training setup of an LLM as a Markov chain among several random variables. Let $\mathcal{K}$ denote a set of random facts, such as a specific list of biographies, drawn from a universe of potential facts. Let the tuple $(\Theta,\mathcal{K})$ denote a “world state” parameterizing the training data generation process. Let $Z$ denote the training data, and $W$ the learned weights, obtained by applying a randomized training algorithm to $Z$. It follows that we have a Markov chain:
@@
We present two arguments: regularization techniques and learning-theoretic pressures that limit pure memorization, and the fact that the total memory capacity is “crowded out” by structured knowledge, leaving little budget for the unstructured random facts.
@@ A.2 Regularization limits memorization
By the data processing inequality, we have $I(W;\mathcal{K},\Theta)\leq I(W;Z)$, and it can be argued that the latter is much smaller than the raw parameter count.
Some of the most established theories of generalization in deep learning include the minimum description length (MDL) principle Rissanen (1978); Gr"unwald (2007); Gr"unwald and Roos (2019), PAC-Bayes McAllester (1999, 2003); Catoni (2007); Alquier (2021), and mutual-information generalization bounds Xu and Raginsky (2017); Bu et al. (2020); Steinke and Zakynthinou (2020). These ideas are interconnected Blum and Langford (2003); Achille and Soatto (2018); Hellstr"om et al. (2025); Grunwald et al. (2021), and they roughly all establish that limited memorization of training data is a key factor in generalization.
@@
These theories posit that the key to explaining generalization lies in controlling memorization, quantified by $I(W;Z)$. This is done through implicit bias imposed by the various regularization mechanisms, such as mini-batch SGD, weight decay, dropout, and early stopping, and the preference of flat minima. If this line of reasoning is correct, then we expect the model to have a tendency to reduce $I(W;Z)$, as long as doing so does not increase the training loss too much.
@@
Suppose that the world is parameterized by random variables $(\Theta,\mathcal{K})$, where $\Theta$ represents structured components (e.g., linguistic syntax, logical rules, universal regularities) and $\mathcal{K}$ represents random facts. We can assume that $\mathcal{K}$ is independent of $\Theta$.
@@
Information theory dictates that these two components compete for the model’s finite capacity. If we assume $\mathcal{K}$ is independent of $\Theta$, then by decomposition of multivariate information Williams and Beer (2010), we have:
@@
Crucially, this is not just a static tradeoff but a dynamic one where $\Theta$ takes precedence. Pan et al. (2025) demonstrate theoretically (via Kolmogorov Structure Functions) and empirically that LLMs exhibit frequency-dependent learning: they prioritize compressing pervasive syntactic structures ($\Theta$) before acquiring rarer factual knowledge ($\mathcal{K}$). This observation is consistent with the two-stage training dynamics of Transformers Gong et al. (2025), as well as the empirical difficulty for trained models to recall high-entropy facts compared to structured knowledge Huang et al. (2025c).

### v2

@@ 1 Introduction
Despite their transformative impact, Large Language Models (LLMs) are hindered by hallucinations—the generation of confident and plausible yet factually incorrect statements. Recently, an influential line of work Kalai et al. (2025); Mohsin et al. (2025); Xu (2025); Wu et al. (2025); Gumaan (2025) provided theoretical explanations of hallucination from a statistical learning perspective, by viewing an LLM as implementing a binary classifier over random facts—facts that are unstructured and cannot be logically inferred unless encountered during training, such as phone numbers or biographical details. By a no-free-lunch-style logic, since generalization is impossible on these facts, the model must make uninformed guesses on unseen random facts, leading to systematic inference errors. In the classification setting, hallucination corresponds to false positives resulting from such guesses.
As a natural workaround, researchers have promoted abstention Kalai et al. (2025); Wen et al. (2025), where an LLM is expected to provide an “I don’t know” response on unseen random factual queries. Theoretically, we can further simplify the task by adopting a “closed-world assumption,” where all unseen potential facts are treated as non-facts at evaluation time  [^1] Alternatively, one can assume all true facts are seen during training.. Under this setting, an ideal LLM can be both informative and non-hallucinating on factual queries, as long as it accurately distinguishes the finitely many known facts from all other (potentially true but unknown) non-facts. This reduces hallucination mitigation to a sample-efficient memorization task.
However, empirical evidence suggests that LLMs struggle to perform this memorization task reliably. Even when permitted to abstain, models continue to generate high-confidence hallucinations and also exhibit “over-refusal” on legitimate queries Cheng et al. (2024); Brahman et al. (2024); Zhu et al. (2025), reflecting a precision-recall trade-off. This points to a different bottleneck beyond no-free-lunch: the finite capacity of model parameters forces lossy compression of the training data, making memorization of specific, unstructured facts non-trivial.
This information-theoretic perspective connects to several recent works Pan et al. (2025); Shi et al. (2025); Mohsin et al. (2025); Kim (2025b, a) that attribute hallucination to distortion incurred when compressing an infinitely complex world of knowledge into a model with finite capacity. Despite these insights, two gaps remain. First, “compression causes errors” does not explain the shape of the errors: while forgetting is natural under limited capacity, it is unclear why hallucination is particularly prevalent. Second, existing compression-based explanations are either high-level informal arguments or assume an infinite number of facts, whereas our closed-world assumption restricts facts to a finite set. This motivates a more rigorous study in a simplified setting:
@@
In this work, we argue that these hallucinations are driven by the asymmetry between facts and non-facts. Specifically, if facts are sparse and randomly distributed in a vast universe of plausible statements, then an LLM with limited memory capacity will tend to accept many non-facts as facts.
To formalize this intuition, we model factuality judgment as a membership testing problem. We view each plausible statement as an element $i$ in a universe $\mathcal{U}$, and the set of known facts as a key set $\mathcal{K}\subseteq\mathcal{U}$. Given a query $i\in\mathcal{U}$, the model outputs a confidence score $\hat{x}_{i}\in[0,1]$ indicating its belief that $i\in\mathcal{K}$; under this view, hallucinations correspond to non-keys being assigned high confidence. The score $\hat{x}_{i}$ can be evaluated under generic error metrics that measure the discrepancy between $\hat{x}_{i}$ and the true label $\mathbbm{1}\{i\in\mathcal{K}\}$. The goal of a membership tester is to use a small memory budget to store the keys and achieve small expected error on queries. This abstraction subsumes approximate membership data structures [^2] We will call these data structures “filters” for convenience. such as Bloom filters Bloom (1970), for which the query output is limited to $\{0,1\}$.
@@ 1.1 Our contributions
A rate-distortion theorem for membership testing. We consider the sparse limit where $\frac{|\mathcal{K}|}{|\mathcal{U}|}\to 0$. In this regime, we identify the exact memory-error trade-off in the form of a rate-distortion theorem. Specifically, the minimum memory budget per key is characterized by the minimum KL-divergence between key and non-key output distributions.
@@
Hallucination as the optimal mode of error. Our theory provides two perspectives on why it is memory-efficient for LLMs to hallucinate on random facts. First, consider a model tasked with estimating the probability $\mathbb{P}[i\in\mathcal{K}]$ that a potential fact $i\in\mathcal{U}$ is true, where the output $\hat{x}_{i}\in[0,1]$ is evaluated with logarithmic or cross-entropy loss, consistent with the maximum likelihood objective and previously proposed training/evaluation methods for factual knowledge Kadavath et al. (2022); Cheng et al. (2024).
The unique loss-minimizing query output under this metric, given a fixed memory capacity, is to assign high confidence to all facts, while assigning either zero or fact-level confidence to non-facts. In other words, the model simultaneously recalls all facts and hallucinates on a fraction of non-facts. We also empirically verify this tendency on synthetic data. For a probability estimation task, hallucination, instead of systematic forgetting or uniform uncertainty, is the natural mode of error.
Classifier via thresholding and two-sided filters. Our second perspective establishes that any LLM decision mechanism based on score thresholding—whether the scores are derived from generative probabilities Kalai et al. (2025) or the probability estimation output above—is subject to the fundamental memory-error trade-off of two-sided filters, a generalization of Bloom-type filters Bloom (1970) allowing both false positives and false negatives. Unlike the first perspective, this result applies to all LLM-based classification mechanisms and does not assume optimality.
@@
Further discussions on the memory capacity. Although modern models have billions of parameters, the effective memory budget for storing a particular family of random facts can be far smaller. Appendix A provides two explanations: (i) modern networks are encouraged to minimize pure memorization (through various regularization and MDL/PAC-Bayes viewpoints), and (ii) structured knowledge (e.g., language and reasoning) and random facts compete for a finite memory budget, where the former takes precedence during learning. We argue that this implicit bias results in a small effective memory budget for random facts, leading to the observed high-confidence hallucinations. Our view resonates with the “memorization is necessary” argument by Feldman et al. Feldman (2020); Feldman and Zhang (2020); Brown et al. (2021): while (implicit) regularization reduces the overfitting to noise, its tendency to limit memorization can be detrimental on the model’s performance on long-tail, high-entropy facts, leading to hallucination.
Conversely, our analysis supports the effectiveness of additional fine-tuning on unstructured random facts, which encourages the model to allocate more memory budget. Our framework also justifies why incorporating external information, such as using RAG Lewis et al. (2020), is effective in mitigating hallucinations: memory budget is no longer a limiting factor when non-parametric memory is present.
@@ 3 A Rate-Distortion Theorem for Membership Testers
Minimizing $F_{p}$ over the feasible sets (distributions $\mu_{K},\mu_{N}$ satisfying error constraints) defines the “rate–distortion” function $R_{p}(\varepsilon_{K},\varepsilon_{N})$. We show that, for each $p$, the “rate” $R_{p}$ is both necessary (Corollary˜3.6) and sufficient (Lemma˜3.7) for achieving the given error level.
@@ 4 Hallucination on Random Facts
Fix a universe $\mathcal{U}$ of $u$ unstructured potential facts; each $i\in\mathcal{U}$ is a plausible natural-language claim (e.g., detailed biographies as in Allen-Zhu and Li (2024)). The set of known facts is the key set $\mathcal{K}\subseteq\mathcal{U}$. In the random-facts regime that motivates our study, $\mathcal{K}$ is a subset of size $n$ drawn uniformly at random from $\mathcal{U}$. This regime intentionally isolates non-generalizable factual knowledge, forcing the LLM to behave like a membership tester for $\mathcal{K}$.
Consider an LLM that is trained (among other tasks) to distinguish facts in $\mathcal{K}$ from all other non-facts. Viewing the LLM as a membership tester for $\mathcal{K}$, Init corresponds to training on labeled data, and Query corresponds to the inference-time response (or internal evaluation) to a plausible claim $i\in\mathcal{U}$. We study two natural regimes of factuality judgment:
@@
    Probability estimation. The LLM generates a confidence score $\hat{x}_{i}\in[0,1]$ as an estimate of $\mathbb{P}[i\in\mathcal{K}]$.
@@ 4.1 Probability estimation variant
Figure˜1 illustrates the shape of the optimal non-fact distribution in blue. The proof appears in Appendix˜C. To solve this optimization problem over probability distributions, we use variational calculus and verify the KKT conditions.
Hallucination channel and its cost. The optimal strategy for a given memory budget is to output a single high-confidence value $x^{*}$ on all facts, while assigning the same high-confidence value $x^{*}$ to a $q^{*}$ fraction of non-facts, forming a “hallucination channel.” These non-facts are indistinguishable from facts by any downstream procedure that observes only $\hat{x}$. Moreover, the per-key memory lower bound simplifies to
@@ 4.2 Binary decision variant
We now consider any binary decision rule on the potential fact universe induced by the LLM. This covers thresholding the probability estimation in Section˜4.1, as well as any classifier arising from generative hallucination probabilities (e.g., in Kalai et al. (2025)). Formally, any such pipeline induces a (randomized) decision $\hat{x}_{i}\in\{0,1\}$ for each $i\in\mathcal{U},$ indicating either acceptance or rejection. We place error constraints on false negative rates (FNR) and false positive rates (FPR):
@@
This memory-error frontier formalizes the trade-off between hallucination and over-refusal observed in practice Cheng et al. (2024); Brahman et al. (2024). Theorem 4.3 is agnostic to how $\hat{x}$ is produced. Whether the decision is derived from a calibrated probability, a raw logit, a log-likelihood ratio, or a multi-stage verification chain that outputs a scalar score, the final thresholded mapping is exactly a two-sided filter with some $(\varepsilon_{K},\varepsilon_{N})$. Consequently, no choice of threshold can beat the memory-error frontier: a conservative approach to factual answers may reduce $\varepsilon_{N}$, but then either the memory budget or $\varepsilon_{K}$ must increase (more forgetting / abstention).
@@
To complete the argument for filters, Appendix D.3 gives an explicit construction of a hash-based two-sided filter that achieves the space lower bound up to $o(1)$ bits per key.
@@ 4.3 Experiments
5 Acknowledgements
We thank Aravindan Vijayaraghavan and Clifford Stein for helpful discussions. Anxin Guo is partially supported by the National Science Foundation through the NSF grants ECCS-2216970 and CCF-2154100 through Aravindan Vijayaraghavan. Jingwei Li is partially supported by NSF grant CCF-2218677 and ONR grant ONR-13533312 through Clifford Stein.
@@ Appendix A Detailed Discussion of the Information Budget for Random Facts
We provide a self-contained argument in support of our modeling choice: namely, that the effective memory budget available for storing random facts is orders of magnitude smaller than the raw parameter count.
@@ A.1 Setup: random facts, training data, and learned weights
We model LLM training via a Markov chain relating several random variables. Let $\mathcal{K}$ denote a set of random facts (e.g., a specific list of biographies) drawn from a large universe of potential facts. Let the tuple $(\Theta,\mathcal{K})$ denote a “world state” that parameterizes the training data generation process. Let $Z$ denote the training data, and let $W$ denote the learned weights obtained by applying a randomized training algorithm to $Z$. This induces the Markov chain:
@@
We present two complementary arguments. First, regularization techniques and learning-theoretic pressures limit pure memorization. Second, much of the model’s effective capacity is “crowded out” by structured knowledge, leaving little budget for unstructured random facts.
@@ A.2 Regularization limits memorization
By the data processing inequality, we have $I(W;\mathcal{K},\Theta)\leq I(W;Z)$; moreover, standard compression- and generalization-based perspectives suggest that $I(W;Z)$ can be far smaller than the raw parameter count.
Some of the most established theories of generalization in deep learning include the minimum description length (MDL) principle Rissanen (1978); Gr"unwald (2007); Gr"unwald and Roos (2019), PAC-Bayes McAllester (1999, 2003); Catoni (2007); Alquier (2021), and mutual-information generalization bounds Xu and Raginsky (2017); Bu et al. (2020); Steinke and Zakynthinou (2020). These ideas are interconnected Blum and Langford (2003); Achille and Soatto (2018); Hellstr"om et al. (2025); Grunwald et al. (2021), and they all roughly establish that limited memorization of training data is a key factor in generalization.
@@
These theories posit that generalization is closely tied to controlling memorization, as quantified by $I(W;Z)$. In practice, $I(W;Z)$ is shaped by the implicit bias of common regularization mechanisms (e.g., mini-batch SGD, weight decay, dropout, and early stopping) and by the preference for flat minima. If this line of reasoning is correct, then we expect training to reduce $I(W;Z)$ whenever doing so does not significantly increase the training loss.
@@
It is also worth noting that a line of work Feldman (2020); Feldman and Zhang (2020); Brown et al. (2021); Feldman et al. (2025) argues that memorization of samples (and even noise) is necessary for generalization in certain cases. Our framework complements this perspective: limiting memorization via (implicit or explicit) regularization, while intended to filter out noise, also inevitably excludes incompressible long-tail random facts. Under an insufficient memory budget, our theory predicts that the model is forced to hallucinate as the information-theoretically optimal response.
@@
Suppose that the world is parameterized by random variables $(\Theta,\mathcal{K})$, where $\Theta$ represents structured components (e.g., linguistic syntax, logical rules, universal regularities) and $\mathcal{K}$ represents random facts. We assume that $\mathcal{K}$ is independent of $\Theta$ (so $I(\mathcal{K};\Theta)=0$).
@@
Information theory dictates that these two components compete for the model’s finite capacity. If we assume $\mathcal{K}$ is independent of $\Theta$, then a standard decomposition of multivariate information Williams and Beer (2010) gives:
@@
Crucially, this is not just a static tradeoff but a dynamic one where $\Theta$ takes precedence. Pan et al. (2025) demonstrate theoretically (via Kolmogorov Structure Functions) and empirically that LLMs exhibit frequency-dependent learning: they prioritize compressing pervasive syntactic structures ($\Theta$) before acquiring rarer factual knowledge ($\mathcal{K}$). This observation is consistent with the two-stage training dynamics of Transformers Gong et al. (2025), as well as the empirical difficulty for trained models to recall high-entropy facts compared to structured knowledge Huang et al. (2025c). Arpit et al. (2017) also concluded that regularized SGD for deep networks would prioritize learning low-entropy “simpler” samples over high-entropy noise.